OpenStudy (anonymous):

use subsitute to solve x-3y=-4 2x+6y=5

OpenStudy (marigirl):

The equation x-3y=-4 make x the subject x=3y-4 Then 'substitute' the x into the second equation 2x+6y=5 \[2(3y-4)+6y=5\] \[6y-8+6y=5\]

OpenStudy (marigirl):

\[12y=13\] \[y=\frac{ 13 }{ 12 }\]

OpenStudy (marigirl):

to find the value of x, substitute y=13/12 into any of the two original equations

OpenStudy (anonymous):

What did I do wrong? I have 12y=11or y=11/12

OpenStudy (marigirl):

What was the x expression that you substituted into the second equation? @Bam13

OpenStudy (anonymous):

I see i added instead of multiplying by a mistake 2(-4+3y) -6+6y oops lol thanks so much

OpenStudy (marigirl):

all good! A medal will be nice :)